# Emergent spin order and steady-state superradiance in one-dimensional baths

### v1

@@ Abstract
Spontaneous collective decay in driven atomic ensembles can generate coherence far from equilibrium, as illustrated by superradiant lasers where decay into a single-mode cavity synchronizes atomic phases into a macroscopic dipole and yields superradiant emission of light with an ultranarrow spectrum. Whether similar ordering persists in multimode reservoirs with propagation and competing collective decay channels remains an open question. We address this problem by analyzing atoms coupled to one-dimensional electromagnetic baths through two models. The first is a ring cavity supporting two bright collective decay channels, and the second is a bidirectional waveguide where, in addition to competition between channels, propagation induces coherent dipole–dipole interactions. For suitable incoherent pumping strengths, the dynamics enters a synchronization window in which collective decay overcomes disordering processes, leading to spontaneous steady-state phase ordering and superradiant emission. We extract the thresholds marking the onset of synchronization and show that the maximum intensity scales as $N^{2}$ in both models. The resulting order is not described by a single macroscopic dipole: in the ring cavity spontaneous chirality emerges at the level of individual trajectories, while the waveguide develops a local chirality with different orders dominating opposite ends of the atomic array. The analysis of the emitted light spectrum reveals a linewidth that seems to narrow with increased system size in the ring cavity, while narrowing in the waveguide remains inconclusive within accessible numerics. These results clarify how competition and propagation shape emergent order in one-dimensional reservoirs and identify regimes where steady-state superradiance may arise beyond the Dicke limit.
@@ I INTRODUCTION
The spontaneous emergence of order in systems composed of many interacting subsystems is pervasive across biology, chemistry, and the social sciences [Haken83]. In physics, disorder–order transitions have been extensively studied in thermodynamic equilibrium, yet some of the most striking manifestations occur far from equilibrium, where a featureless external energy input can stabilize robust ordered steady states and generate macroscopic coherence. A prime example is the conventional laser [Haken83, Haken84, Mandel95], where atoms in a “high-Q” single-mode cavity pumped above the lasing threshold, generate coherent light via stimulated emission. The intracavity field both quantifies the degree of organization and provides the feedback that enforces it.
@@
Another example of spontaneous order in the form of macroscopic quantum coherence is that of a superradiant laser, realized in the “bad-cavity” limit where photons leak out faster than the time needed for stimulated emission [Meiser09, Meiser10, Zhang18, Liu20, Reilly25]. Here, the cavity acts as a shared electromagnetic reservoir that mediates all-to-all interactions between the atoms, and endows the system with permutational symmetry if the atoms are placed at the cavity antinodes. As a result, the atoms spontaneously decay via a single bright jump operator [Dicke54]. For suitable pumping strengths, the atoms relax to a superradiant state [Dicke54], emitting a field whose intensity scales quadratically with atom number and whose linewidth can fall below that of a single-atom, suppressing noise from dephasing or decay outside of the cavity. The repeated action of the single jump operator synchronizes the atomic phases [Acebron05, Xu13, Zhu15], establishing a macroscopic dipole that is resilient against quantum fluctuations. In contrast to a conventional laser, phase information is thus stored primarily in the atomic ensemble rather than in the cavity field.
@@
An open question is whether steady-state spontaneous spin order – together with superradiant light emission and spectral narrowing – can be stabilized away from the previously described “Dicke limit”. If atoms decay into a multimode electromagnetic bath with continuous spectrum, the competition from the resulting collective jump operators – together with light-propagation effects encoded in the Hamiltonian – may preclude the emergence of an ordered phase. Nevertheless, prior work [Zhu15, Bychek25] indicates that the Dicke limit is not required to observe order-like signatures and line narrowing. Furthermore, recent work on transient self-organization in extended systems – including superradiant bursts [Masson20, Masson22, Sierra22, robicheaux2021theoretical, Liedl24, Rubies23], directional photon emission [Dicke54, Cardenas23, Lukin25], and phase ordering detected via heterodyne measurements [Zhang25B] – is encouraging. Yet a clear physical picture of how competition and propagation shape the emergent order remains elusive, as does the question of whether linewidths below the natural single-atom value and steady-state superlinear (i.e., superradiant) emission are feasible beyond the Dicke limit.
@@
Here, we address these questions by elucidating the roles of multimode competition and light propagation in the steady-state emission of atoms coupled to one-dimensional (1D) reservoirs. We consider different spin models that progressively incorporate these effects: (i) a ring cavity [Lee25], which supports two bright jump operators, each driving competing orders for the atomic coherence and emitted light, and (ii) a bidirectional waveguide, where in addition to competition between atomic collective modes, light propagation manifests through coherent dipole–dipole interactions mediated by the waveguide. Beyond serving as minimal models to explore competition and propagation, these scenarios are directly relevant to experimental platforms such as cold atoms coupled to microring resonators [Zhou24, Suresh25, Zhou25], ring and bow-tie cavities [Cline25, Simon20, Chen22] and nanofibers [Vetsch10, Goban12, Gouraud15, Solano17].
@@ II.1 Spin models
We consider a system of $N$ two-level atoms, each with a resonance frequency $\omega_{0}$, interacting with a common electromagnetic reservoir at an individual decay rate $\Gamma_{\rm 1D}$. Atoms also undergo parasitic individual decay at rate $\Gamma^{\prime}$ and are incoherently pumped at a rate $w$. After tracing out the electromagnetic field, the atomic dynamics under the Born-Markov approximation are governed by the master equation [Gruner96, Dung02, Asenjo17]:
@@
where
$\hat{H}$ $=\hbar\sum_{n,m=1}^{N}J_{nm}\hat{\sigma}_{+}^{n}\hat{\sigma}_{-}^{m},$ (2a)
@@
and the pumping term in the Lindbladian, $\mathcal{L}_{w}[\hat{\rho}$], is obtained from $\mathcal{L}_{\Gamma}[\hat{\rho}$] by swapping $\hat{\sigma}_{+}^{n}\leftrightarrow\hat{\sigma}_{-}^{n}$. Here, $\hat{\sigma}_{+}^{n}\equiv\mathinner{|{e_{n}}\rangle}\mathinner{\langle{g_{n}}|}$ denotes the coherence operator for emitter $n$. The coefficients $\{J_{nm},\Gamma_{nm}\}$ encode the coherent and dissipative interactions, respectively, between atoms $n$ and $m$. These interactions are mediated by the shared environment and are fully determined by the electromagnetic field propagator (i.e., the electromagnetic Green’s function) between pairs of atoms, evaluated at the resonance frequency [Asenjo17].
@@
where $\hat{\mathcal{O}}_{\nu}=\sum_{n=1}^{N}b_{\nu}^{n}\hat{\sigma}_{-}^{n}$ are collective jump operators with rates $\Gamma_{\nu}$ (and with spatial profile given by the eigenvectors $\bm{b}_{\nu}$ of $\bm{\Gamma}$). These modes may be superradiant ($\Gamma_{\nu}>\Gamma_{\rm 1D}$) or subradiant ($\Gamma_{\nu}<\Gamma_{\rm 1D}$). Alternatively, $\mathcal{L}_{\text{dis}}[\hat{\rho}$] can be rewritten in terms of directional jump operators [Carmichael2000, Clemens2003].
The steady state of the density matrix dynamics described by Eq. (1) is unique. This follows from Evans’ theorem [Evans77, Buca12], which states that the steady state is unique whenever the dynamical generators of the master equation – the Hamiltonian and all jump operators – span the full operator space. In our case, this condition is met due to the presence of individual parasitic decay and incoherent pumping processes.
Once the atomic evolution described by Eq. (1) is obtained, the properties of the emitted electromagnetic field follow from the input-output relations [Lalumiere13, Caneva15, Xu15_2], which express the field in terms of atomic operators and the electromagnetic field propagator. The photon emission rate, which is proportional to the field intensity, reads
@@
This observable reveals signatures of collective dynamics. Examples include transient superradiant bursts when the atoms decay from a fully inverted state [Dicke54], or suppressed emission when subradiant states are populated [Albrecht19]. Collective behavior manifests itself not only in the total emitted intensity, but also in the spectrum of the emitted light, defined for a field operator $\hat{E}$ as
$S_{\hat{E}}(\omega)\equiv 2\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{E}^{\dagger}(t_{ss}+\tau)\hat{E}(t_{ss})\rangle\right],$ (5)
where $t_{ss}$ denotes the time at which the steady state has been reached.
@@
with collective decay rate $N\Gamma_{\rm 1D}$. The positive frequency component of the field operator is directly related to the jump operator and reads [Bonifacio71]
@@
This model, depicted in Fig. 1(a), is realized by atoms resonantly coupled to a bad cavity, i.e., a cavity whose photon leakage rate far exceeds any other timescale [Bonifacio71, Meiser10]. This is the conventional model of Dicke superradiance [Dicke54] and is, to date, the only one where superradiant lasing has been demonstrated to exist [Meiser09, Meiser10, Bohnet12, Kristensen23].
Ring cavity (two collective jump operators): For this model, we again set the Hamiltonian to zero but now introduce position-dependent dissipative couplings [Hummer13, Lee25]:
@@
where $k$ is the wavenumber of the cavity mode and $z_{n}$ are the atom positions. This yields two bright collective jump operators corresponding to left (counterclockwise) and right (clockwise) emission [Pichler15]
@@
This model, depicted in Fig. 1(b), is physically realized by resonantly coupling atoms to a ring cavity and is designed to probe mode competition. Using the input-output relation [Caneva15] together with the Green’s function for a ring cavity [Hummer13, Lee25], the left- and right- propagating fields at any position $z$ are
@@
Bidirectional waveguide (two jump operators and Hamiltonian evolution): Finally, this model incorporates both Hamiltonian evolution and position-dependent dissipation. The interaction coefficients are given by [Asenjo17, Pichler15]:
$J_{nm}-{\rm i}\frac{\Gamma_{nm}}{2}=-{\rm i}\frac{\Gamma_{\rm 1D}}{2}e^{{\rm i}k|z_{n}-z_{m}|},$ (12)
@@
For both the ring cavity and the waveguide, we assume that atoms form ordered arrays with spacing $d$, i.e., $z_{n}=nd$, although as we will detail later, the results are expected to extend to disordered arrays. In the “mirror” configuration ($kd=n\pi$ with $n\in\mathbb{Z}$), the coherent part $J_{nm}$ vanishes, and the system supports a single bright mode, reducing the dynamics to that of a single-mode cavity [Chang12].
@@ II.2 Analytical and computational methods
Ordered phases of matter are emergent phenomena that become more robust and well-defined as $N$ grows, which makes the computation of their physical properties challenging away from thermal equilibrium. Equation (1) describes an open quantum system with an exponentially large Hilbert space, rendering exact naive simulations intractable at large $N$. When full permutation symmetry is present (as in a single-mode cavity) the dynamics can be computed exactly with computational cost $\sim N^{3}$ [Xu13]. This can be extended to ring cavities by exploiting partial permutation symmetry in commensurate “superspin” arrays [Lee25]. In particular, for spacings $d=m\lambda_{0}/(2p)$ with $m,p\in\mathbb{Z}$, the ensemble decomposes into $p$ symmetric groups $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$), each represented by a collective spin
@@
where $N_{\alpha}$ is the number of atoms in $\mathcal{S}_{\alpha}$ and the coefficients $\{c_{n,\alpha}\}$ ensure that coherences of atoms within the same symmetric group add in phase (see Appendix A). This reduces the computational complexity to $\sim N^{3p}$ and enables a compact description fully in terms of the collective operators in Eqs. (14). The case $p=1$ reduces to the single macroscopic dipole describing a superradiant laser [Xu13]. Advancing beyond the single-mode analysis of Ref. [Meiser10], we employ this formalism in Sec. IV to characterize the steady-state emission of the ring cavity.
In the waveguide, where both full and partial permutation symmetry are absent, we resort to approximate methods. The simplest is mean field, which neglects correlations by factorizing the density matrix and treating each emitter as a classical dipole. To include leading quantum corrections, we employ the Truncated Wigner Approximation (TWA) for dissipative spin systems, which yields $2N$ stochastic differential equations, where correlated white noise terms model quantum fluctuations [Polkovnikov10, Schachenmayer15, Mink22, Mink23, Guardiola25]. The reliability of this approach is supported by benchmarks against exact small-$N$ simulations, together with recent evidence that few-body observables in superradiant decay with all-to-all coupling are well captured by mean-field trajectories under homodyne unraveling [Zhang25]. We use this method throughout the manuscript to assess the impact of quantum fluctuations.
@@ II.3 Mean field and classical synchronization
Employing a mean-field approximation provides analytical insight into the emergent spin order by interpreting it as phase synchronization. Here, we follow Ref. [Zhu15] and present mean-field equations that we will later apply to our particular spin models. In the mean-field approach, quantum correlations are neglected by assuming a factorized density matrix, $\rho(t)\approx\bigotimes_{n=1}^{N}\rho_{n}(t)$. Each atom is then described by an inversion, coherence, and a phase, $\{s_{n}^{z},s_{n}^{\perp},\phi_{n}\}$, defined as
@@
which reduce Eq. (1) to the mean-field equations [Zhu15]
@@
where $\phi_{n\ell}\equiv\phi_{n}-\phi_{\ell}$. Relative phases determine the coherent and collective dissipative evolution, as seen from the terms proportional to $J_{nm}$ and $\Gamma_{nm}$.
Within mean field, phase dynamics map to a Sakaguchi–Kuramoto model [Zhu15], a standard description of synchronization and pattern formation in networks of classical limit cycles [Acebron05]. The basic Kuramoto model for phases $\{\phi_{\ell}\}$ assumes permutational symmetric coupling and reads [Kuramoto84, Strogatz00, Acebron05]
@@
showing that phases are attracted toward the order parameter phase $\psi$, which causes the growth of $r$, increasing attraction and establishing a positive feedback loop that stabilizes at nonzero $r$. Cosine-type phase interaction (like the ones describing the coherent interactions in the phase mean-field equation) give instead an integrable model [Watanabe93]: its evolution retains memory of the initial conditions, and disordered configurations do not spontaneously synchronize. When sine- and cosine-type all-to-all couplings compete, the balance between the two dictates whether the system synchronizes or disorder persists [Panaggio15]. Extensions beyond permutation-symmetric coupling can yield reduced synchronization or chimera states, where synchronized and unsynchronized oscillators coexist [Sakaguchi86, Kuramoto02, Abrams04, Panaggio15].
Even without coherent interactions, Eq. (16c) yields richer behavior than the standard Kuramoto model [Zhu15]. Effective couplings are inhomogeneous and time dependent through $s_{\ell}^{z}$ and $s_{\ell}^{\perp}$. Positive steady-state couplings require $s_{\ell}^{z}>0$ and $s^{\perp}_{\ell}\neq 0$, which define a finite synchronization window for $w$. When $w\in[w^{(l)},w^{(u)}]$, the couplings are positive at long times and phase synchronization is possible. The boundaries of the synchronization window are set by $\{J_{nm},\Gamma_{nm}\}$.

### v2

@@ Abstract
Spontaneous collective decay in incoherently driven atomic ensembles generates macroscopic coherence in the steady state, as exemplified by superradiant lasing in single-mode cavities. Whether spontaneous order persists in multimode reservoirs with competing collective decay channels and light propagation remains an open question. We address this problem by analyzing atoms coupled to one-dimensional electromagnetic baths through two models: a ring cavity with two bright decay channels, and a bidirectional waveguide where, in addition to competition between channels, propagation induces Hamiltonian dipole–dipole interactions. For suitable pumping strengths, both models enter a synchronization window leading to steady-state phase ordering and superradiant emission, scaling as $N^{2}$. The resulting order is not described by a single macroscopic dipole: in the ring cavity spontaneous chirality emerges at the level of individual trajectories, while the waveguide develops a local chirality with different orders dominating opposite ends of the array. The spectral analysis of the emitted light seems to indicate line narrowing with increased system size in the ring cavity. Line narrowing in the waveguide remains inconclusive within accessible numerics. Our results demonstrate how competition and propagation shape emergent order beyond the Dicke limit.
@@ I INTRODUCTION
The spontaneous emergence of order in systems composed of many interacting subsystems is pervasive across biology, chemistry, and the social sciences [27]. In physics, disorder–order transitions have been extensively studied in thermodynamic equilibrium, yet some of the most striking manifestations occur far from equilibrium, where a featureless external energy input can stabilize robust ordered steady states and generate macroscopic coherence. A prime example is the conventional laser [27, 28, 38], where atoms in a “high-Q” single-mode cavity pumped above the lasing threshold, generate coherent light via stimulated emission. The intracavity field both quantifies the degree of organization and provides the feedback that enforces it.
@@
Another example of spontaneous order in the form of macroscopic quantum coherence is that of a superradiant laser, realized in the “bad-cavity” limit where photons leak out faster than the time needed for stimulated emission [42, 41, 73, 36, 50]. Here, the cavity acts as a shared electromagnetic reservoir that mediates all-to-all interactions between the atoms, and endows the system with permutational symmetry if the atoms are placed at the cavity antinodes. As a result, the atoms spontaneously decay via a single bright jump operator [18]. For suitable pumping strengths, the atoms relax to a superradiant state [18], emitting a field whose intensity scales quadratically with atom number and whose linewidth can fall below that of a single-atom, suppressing noise from dephasing or decay outside of the cavity. The repeated action of the single jump operator synchronizes the atomic phases [2, 66, 76], establishing a macroscopic dipole that is resilient against quantum fluctuations. In contrast to a conventional laser, phase information is thus stored primarily in the atomic ensemble rather than in the cavity field.
@@
An open question is whether steady-state spontaneous spin order – together with superradiant light emission and spectral narrowing – can be stabilized away from the previously described “Dicke limit”. If atoms decay into a multimode electromagnetic bath with continuous spectrum, the competition from the resulting collective jump operators – together with light-propagation effects encoded in the Hamiltonian – may preclude the emergence of an ordered phase. Nevertheless, prior work [76, 9] indicates that the Dicke limit is not required to observe order-like signatures and line narrowing. Furthermore, recent work on transient self-organization in extended systems – including superradiant bursts [40, 39, 57, 51, 35, 52], directional photon emission [18, 11, 37], and phase ordering detected via heterodyne measurements [71] – is encouraging. Yet a clear physical picture of how competition and propagation shape the emergent order remains elusive, as does the question of whether linewidths below the natural single-atom value and steady-state superlinear (i.e., superradiant) emission are feasible beyond the Dicke limit.
@@
Here, we address these questions by elucidating the roles of multimode competition and light propagation in the steady-state emission of atoms coupled to one-dimensional (1D) reservoirs. We consider two spin models that progressively incorporate these effects: (i) a ring cavity [34], which supports two bright jump operators, each driving competing orders for the atomic coherence and emitted light, and (ii) a bidirectional waveguide, where in addition to competition between atomic collective modes, light propagation manifests through coherent dipole–dipole interactions mediated by the waveguide. Beyond serving as minimal models to explore competition and propagation, these scenarios are directly relevant to experimental platforms such as cold atoms coupled to microring resonators [75, 60, 74], ring and bow-tie cavities [17, 15, 14] and nanofibers [62, 22, 24, 58].
@@ II.1 Spin models
We consider a system of $N$ two-level atoms, each with a resonance frequency $\omega_{0}$, interacting with a common electromagnetic reservoir at an individual decay rate $\Gamma_{\rm 1D}$. Atoms also undergo parasitic individual decay at rate $\Gamma^{\prime}$ and are incoherently pumped at a rate $w$. After tracing out the electromagnetic field, the atomic dynamics under the Born-Markov approximation are governed by the master equation [25, 20, 5]:
@@
where (making $\hbar=1$)
$\hat{H}$ $=\sum_{n,m=1}^{N}J_{nm}\hat{\sigma}_{+}^{n}\hat{\sigma}_{-}^{m},$ (2a)
@@
and the pumping term in the Lindbladian, $\mathcal{L}_{w}[\hat{\rho}$], is obtained from $\mathcal{L}_{\Gamma}[\hat{\rho}$] by swapping $\hat{\sigma}_{+}^{n}\leftrightarrow\hat{\sigma}_{-}^{n}$. Here, $\hat{\sigma}_{+}^{n}\equiv\mathinner{|{e_{n}}\rangle}\mathinner{\langle{g_{n}}|}$ denotes the coherence operator for emitter $n$. The coefficients $\{J_{nm},\Gamma_{nm}\}$ encode the coherent and dissipative interactions, respectively, between atoms $n$ and $m$. These interactions are mediated by the shared environment and are fully determined by the electromagnetic field propagator (i.e., the electromagnetic Green’s function) between pairs of atoms, evaluated at the resonance frequency [5].
@@
where $\hat{\mathcal{O}}_{\nu}=\sum_{n=1}^{N}b_{\nu}^{n}\hat{\sigma}_{-}^{n}$ are collective jump operators with rates $\Gamma_{\nu}$ (and with spatial profile given by the eigenvectors $\bm{b}_{\nu}$ of $\bm{\Gamma}$). These modes may be superradiant ($\Gamma_{\nu}>\Gamma_{\rm 1D}$) or subradiant ($\Gamma_{\nu}<\Gamma_{\rm 1D}$). Alternatively, $\mathcal{L}_{\text{dis}}[\hat{\rho}$] can be rewritten in terms of directional jump operators [12, 16].
The steady state of the density matrix dynamics described by Eq. (1) is unique. This follows from Evans’ theorem [21, 8], which states that the steady state is unique whenever the dynamical generators of the master equation – the Hamiltonian and all jump operators – span the full operator space. In our case, this condition is met due to the presence of individual parasitic decay and incoherent pumping processes. Furthermore, notice that the choice of incoherent pumping endows Eq. (1) with global $U(1)$ symmetry, that is, invariance under the transformation $\hat{\sigma}_{+}^{n}\rightarrow e^{i\theta}\hat{\sigma}_{+}^{n}$ $\forall$ $n$. As a consequence of these two observations, and in contrast to the case of coherent pumping, the expectation values of all emitter coherences, and hence all phase information, vanish identically in the steady state for any choice of initial state. As we show throughout this work, this information can nonetheless be recovered by examining either single realizations within the ensemble of quantum trajectories or correlation functions of spin operators.
Signatures of order can also be observed at the ensemble-averaged level in the correlation functions of the emitted electromagnetic field. These can be obtained from the atomic evolution via the input-output relations [33, 10, 69], which express the field in terms of atomic operators and the electromagnetic field propagator. The rate of photon emission into the common reservoir, which is proportional to the field intensity, reads
@@
Examples of order manifestations include transient superradiant bursts when the atoms decay from a fully inverted state [18], or suppressed emission when subradiant states are populated [4]. Collective behavior manifests itself not only in the total emitted intensity, but also in the spectrum of the emitted light, defined for a field operator $\hat{E}^{+}$ as
$S_{\hat{E}}(\omega)\equiv 2\,\text{Re}\!\left[\int_{0}^{\infty}d\tau\,e^{-{\rm i}\omega\tau}\langle\hat{E}^{-}(t_{ss}+\tau)\hat{E}^{+}(t_{ss})\rangle\right],$ (5)
[…]
@@
with collective decay rate $N\Gamma_{\rm 1D}$. The positive frequency component of the field operator is directly related to the jump operator and reads [7]
@@
This model, depicted in Fig. 1(a), is realized by atoms resonantly coupled to a bad cavity, i.e., a cavity whose photon leakage rate far exceeds any other timescale [7, 41]. This is the conventional model of Dicke superradiance [18] and is, to date, the only one where superradiant lasing has been demonstrated to exist [42, 41, 6, 30].
Ring cavity (two collective jump operators): For this model, we again set the Hamiltonian to zero but now introduce position-dependent dissipative couplings [29, 34]:
@@
where $k$ is the wavenumber of the cavity mode and $z_{n}$ are the atom positions. This yields two bright collective jump operators corresponding to left (counterclockwise) and right (clockwise) emission [47]
@@
This model, depicted in Fig. 1(b), is physically realized by resonantly coupling atoms to a ring cavity and is designed to probe mode competition. Using the input-output relation [10] together with the Green’s function for a ring cavity [29, 34], the left- and right- propagating fields at any position $z$ are
@@
Bidirectional waveguide (two jump operators and Hamiltonian evolution): Finally, this model incorporates both Hamiltonian evolution and position-dependent dissipation. The interaction coefficients are given by [5, 47]:
$J_{nm}-{\rm i}\frac{\Gamma_{nm}}{2}=-{\rm i}\frac{\Gamma_{\rm 1D}}{2}e^{{\rm i}k|z_{n}-z_{m}|}.$ (12)
@@
For both the ring cavity and the waveguide, we assume that atoms form ordered arrays with spacing $d$, i.e., $z_{n}=nd$, although as we will detail later, the results are expected to extend to disordered arrays. In the “mirror” configuration ($kd=n\pi$ with $n\in\mathbb{Z}$), the coherent part $J_{nm}$ vanishes, and the system supports a single bright mode, reducing the dynamics to that of a single-mode cavity [13].
@@ II.2 Analytical and computational methods
Ordered phases of matter are emergent phenomena that become more robust and well-defined as $N$ grows, which makes the computation of their physical properties challenging away from thermal equilibrium. Equation (1) describes an open quantum system with an exponentially large Hilbert space, rendering exact naive simulations intractable at large $N$. When full permutation symmetry is present (as in a single-mode cavity) the dynamics can be computed exactly with computational cost $\sim N^{3}$ [67]. This can be extended to ring cavities by exploiting partial permutation symmetry in commensurate “superspin” arrays [34]. In particular, for spacings $d=m\lambda_{0}/(2p)$ with $m,p\in\mathbb{Z}$, the ensemble decomposes into $p$ symmetric groups $\mathcal{S}_{\alpha}$ ($\alpha=1,\dots,p$), each represented by a collective spin
@@
where $N_{\alpha}$ is the number of atoms in $\mathcal{S}_{\alpha}$ and the coefficients $\{c_{n,\alpha}\}$ ensure that coherences of atoms within the same symmetric group add in phase (see Appendix A). This reduces the computational complexity to $\sim N^{3p}$ and enables a compact description fully in terms of the collective operators in Eqs. (14). The case $p=1$ reduces to the single macroscopic dipole describing a superradiant laser [67]. Advancing beyond the single-mode analysis of Ref. [41], we employ this formalism in Sec. IV to characterize the steady-state emission of the ring cavity.
In the waveguide, where both full and partial permutation symmetry are absent, we resort to approximate methods. The simplest is mean field, which neglects correlations by factorizing the density matrix and treating each emitter as a classical dipole. To include leading quantum corrections, we employ the Truncated Wigner Approximation (TWA) for dissipative spin systems, which yields $2N$ stochastic differential equations, where correlated white noise terms model quantum fluctuations [49, 55, 44, 43, 26]. The reliability of this approach is supported by benchmarks against exact small-$N$ simulations, together with recent evidence that few-body observables in superradiant decay with all-to-all coupling are well captured by mean-field trajectories under heterodyne unraveling [72]. We use this method throughout the manuscript to assess the impact of quantum fluctuations.
@@ II.3 Mean field and classical synchronization
Employing a mean-field approximation provides analytical insight into the emergent spin order by interpreting it as phase synchronization. Here, we follow Ref. [76] and present mean-field equations that we will later apply to our particular spin models. In the mean-field approach, quantum correlations are neglected by assuming a factorized density matrix, $\rho(t)\approx\bigotimes_{n=1}^{N}\rho_{n}(t)$. Each atom is then described by an inversion, coherence, and a phase, $\{s_{n}^{z},s_{n}^{\perp},\phi_{n}\}$, defined as
@@
which reduce Eq. (1) to the mean-field equations [76]
@@
where $\phi_{n\ell}\equiv\phi_{n}-\phi_{\ell}$. Relative phases determine the coherent and collective dissipative evolution, as seen from the terms proportional to $J_{nm}$ and $\Gamma_{nm}$. In all mean-field simulations, we take the atoms to be fully excited at initial time and incorporate fluctuations by sampling initial conditions from the distribution $p(s^{z},s^{\perp},\phi)\propto e^{-2{s^{\perp}}^{2}}\delta\left(\sqrt{{\left(s^{\perp}\right)}^{2}+{\left(s^{z}\right)}^{2}}-\tfrac{1}{2}\right)$ [55].
Within mean field, phase dynamics map to a Sakaguchi–Kuramoto model [76], a standard description of synchronization and pattern formation in networks of classical limit cycles [2]. The basic Kuramoto model for phases $\{\phi_{\ell}\}$ assumes permutational symmetric coupling and reads [32, 59, 2]
@@
showing that phases are attracted toward the order parameter phase $\psi$, which causes the growth of $r$, increasing attraction and establishing a positive feedback loop that stabilizes at nonzero $r$. Cosine-type phase interaction (like the ones describing the coherent interactions in the phase mean-field equation) give instead an integrable model [63]: its evolution retains memory of the initial conditions, and disordered configurations do not spontaneously synchronize. When sine- and cosine-type all-to-all couplings compete, the balance between the two dictates whether the system synchronizes or disorder persists [46]. Extensions beyond permutation-symmetric coupling can yield reduced synchronization or chimera states, where synchronized and unsynchronized oscillators coexist [53, 31, 1, 46].
Even without coherent interactions, Eq. (16c) yields richer behavior than the standard Kuramoto model [76]. Effective couplings are inhomogeneous and time dependent through $s_{\ell}^{z}$ and $s_{\ell}^{\perp}$. Positive steady-state couplings require $s_{\ell}^{z}>0$ and $s^{\perp}_{\ell}\neq 0$, which define a finite synchronization window for $w$. When $w\in[w^{(l)},w^{(u)}]$, the couplings are positive at long times and phase synchronization is possible. The boundaries of the synchronization window are set by $\{J_{nm},\Gamma_{nm}\}$.